# Particle Trajectory Representation Learning with Masked Point Modeling

### v2

@@ 6 Discussion
The high $F_{1}$ scores achieved for track (0.994) and shower (0.977) token classification validate the efficacy of our SSL approach, while the semantic segmentation results show a successful transfer to dense classification. These metrics are on par with state-of-the-art supervised frameworks, such as SPINE (Drielsma et al., 2021), despite not relying on any labeled examples. This demonstrates that SSL frameworks can robustly capture the complex geometries and energy depositions characteristic of TPC data. However, the suboptimal modeling of the highly-infrequent particles, such as Michel electrons and delta rays, indicates a limitation in our current framework. Future studies should try to remedy this issue, by perhaps reducing the group radius.
Despite these successes, several avenues for improvement remain. The computational overhead introduced by C-NMS could be further optimized to enable real-time processing of large datasets. Exploring alternative SSL paradigms, such as contrastive learning and/or student-teacher distillation frameworks, may also offer benefits. In any case, we hope that the introduction of the PILArNet-M dataset into the community will spur additional research in this domain.
@@
Throughout the development of this work, many changes were introduced to enforce the learning of better representations. We expand on several modifications we tried, but did not result in any meaningful change in token semantics, in Appendix F.
@@ Acknowledgements
Impact Statement
This work advances self-supervised learning (SSL) for 3D particle trajectory analysis in Time Projection Chambers (TPCs), a critical tool in high-energy physics (HEP). By designing SSL techniques that operate on sparse, irregular, and high-precision scientific data, we demonstrate how representation learning can be adapted to the unique constraints of experimental physics.
Beyond particle physics, this work highlights the kinds of innovations needed to make popular machine learning methods viable in hyper-constrained scientific domains. Many experimental sciences – from astrophysics to materials science and biomedical imaging – grapple with large-scale, highly structured, and often non-Euclidean data that traditional deep learning architectures are ill-equipped to handle. Our contributions, such as resolution-agnostic volumetric tokenization and energy-aware masked modeling, provide an example for how SSL can be molded to function in environments that demand both extreme data efficiency and precise physical interpretability.
We hope this work inspires applications of SSL in other experimental sciences that generate high-volume, high-dimensional data but lack sufficient human-labeled annotations. By reducing reliance on manual supervision, improving data efficiency, and promoting open data sharing through PILArNet-M, we aim to accelerate scientific discovery beyond particle physics, helping to bridge the gap between machine learning advances and real-world experimental constraints.
From an ethical standpoint, this research does not introduce new societal risks, as it is confined to fundamental physics research. Its consequences of failure are equivalent to those of traditional reconstruction methods – potential inaccuracies in scientific interpretations – without any direct impact on broader society. Nonetheless, we emphasize the importance of continued scrutiny in evaluating biases and uncertainties inherent to any machine learning model used in scientific inference.
@@ Appendix A Dataset
The PILArNet-M dataset contains 5.2B points containing energy deposited from 28,251,859 individual particle trajectories, over 1,210,080 individual events. As described §3 of the main text, each voxel contains information corresponding to not just the energy deposited, but also a fragment ID, group ID, interaction ID, and semantic type. Though we just use the semantic type in this work, we foresee the other IDs being quite useful for other tasks, like, i.e., latent clustering for individual particle (i.e., group ID) identification. Dataset-wide statistics are shown in Figure 8. We also offer visualizations of individual events, where voxels with the same colors correspond to sharing different IDs, in Figure 9. See (Adams et al., 2020) for detailed information about this dataset.
Figure 8: PILArNet-M statistics. Starting from top left and going clockwise: cluster frequencies, i.e., how prevalent each particle type is on a cluster (trajectory) basis; voxel frequencies, i.e., the total semantic ID makeup of all voxels in the dataset; voxel energy distributions, i.e., a histogram of the energies of any occupied voxel across all events in the dataset. The number in the bar plots represent the raw number of each quantity found in the dataset, while the bars themselves are normalized to sum to 1. Note that in the energy deposition distribution, energy values below 0.001 MeV are not present.
@@
Figure 9: Visual depictions of 7 PILArNet-M events. For each event, voxels are colored according to which integer identifiers they share, for each label type (fragment / cluster ID, group ID, interaction ID, and semantic ID). For ease of visualization, we remove low energy deposits, which would greatly degrade the ability to visualize what each identifier means.
@@ B.1 Comparison of Grouping Strategies
B.1.1 Evaluation Metrics
@@ 2. Duplicated Points Ratio
Prior work mitigates leakage by using large masking ratios ($>$60%), but this only reduces overlap probability rather than eliminating it. For particle trajectory datasets—where spatial correlations are strong—even minor overlaps risk exposing the model to trivial reconstruction shortcuts. Our C-NMS strategy directly optimizes the Pareto frontier of $M$ vs. $D$ (see Figure 3 in the main text), ensuring groups are both complete (low missed points) and isolated (low duplication). This mirrors the non-overlapping structure of successful 2D image masking while adapting to the irregular geometry of particle clouds.
B.1.2 Comparison
Figure 10: Detailed grouping hyperparameter sweep. To find the pareto frontier of grouping methodologies, we sweep over an array of reasonable grouping parameters for each method, reporting the missed ratio $M$ and duplicated ratio $D$ heatmaps for each hyperparameter pair. Clearly, the naïve based approaches result in an extreme amount of duplicated points in any configuration, while the C-NMS-based grouping method empirically guarantees much smaller amounts of each quantity.
Pareto Frontier of Grouping Methodologies.
To systematically compare grouping strategies, we perform parameter sweeps for each method. Initial group centers are sampled via FPS (2048 per event). These parameter ranges are described in Table 3.
@@ B.1.2 Comparison
For FPS-based methods, the number of samples per event are defined as $\lceil|\mathcal{X}|/\alpha\rceil$ where $|\mathcal{X}|$ is the number of points in the event. Sweeps generate $80-220$ configurations per method, with Pareto frontiers computed from the missed/duplicated ratio metrics defined in §B.1.1. Figure 10 provides array of figures showing results for each metric across these sweeps.
@@ Appendix C Hyperparameters
We provide detailed hyperparameter information in the following sections. The full implementation of these models, as well as training code, can be found at https://github.com/DeepLearnPhysics/PoLAr-MAE/.
@@ C.1 Grouping strategy
We chose a group radius equal to 5 voxels (15 mm) such that a 1m track will contain  67 mm, and because it is approximately the characteristic distance of trajectory morphologies. To find the best value for the overlap factor $f$, we run a sweep over values of $f$, and looking at the missed points ratio $M$ and duplicated points ratio $D$ as defined in Appendix B.1.1. Fig. 11 shows this trade-off for different values of $f$. The value of $f$ that jointly minimizes both these scores if $f=0.73,$ and hence we use this value when pretraining with this grouping strategy.
Figure 11: C-NMS overlap parameter $f$ sweep. We find the optimal $f$ parameter for our group radius of 5 voxels by minimizing both the missed ratio $M$ and duplicated ratio $D$. We find $f=0.73$ to be approximately optimal.
@@ Appendix D Pretraining Results
In Figure 13, we show additional visualizations of representations learned by both PointMAE and PoLAr-MAE.
@@ Appendix E Segmentation Results
In Figure 14, we show additional examples of semantic segmentation results for both PointMAE and PoLAr-MAE fine-tuned models. In Table 6, we provide additional evaluation metrics for the fine-tuned PointMAE and PoLAr-MAE models. We also supply confusion matrices for each model, normalized across predictions, in Figure 15.
@@
Figure 14: Further examples of dense classification. Model types are abbreviated with PL meaning PoLAr-MAE, PMAE meaning PointMAE, PEFT meaning parameter-efficient fine-tuning, and FFT meaning full fine-tuning. Best viewed zoomed in.
@@
PointMAE PEFT 0.772 0.965 0.983 0.569 0.572 0.702 0.990 0.998 0.413 0.406 0.948 0.941 0.968 0.919 0.967
PointMAE FFT 0.831 0.963 0.994 0.561 0.807 0.771 0.996 0.997 0.395 0.697 0.962 0.932 0.990 0.969 0.957
@@
Figure 15: Semantic segmentation confusion matrices. Rows denote fine-tuning strategies: PEFT and FFT. Columns denote the pre-trained model used: PointMAE and PoLAr-MAE. Each column sums to 1 such that the diagonal values represent precision metrics for each class.

### v3

@@ 6 Future Work and Discussion
The high $F_{1}$ scores achieved for track (0.994) and shower (0.977) token classification with linear SVMs validate the efficacy of our SSL approach, while the semantic segmentation results show a successful transfer to dense classification. These metrics are on par with state-of-the-art supervised frameworks, such as the Sparse UResNet trained on $>100,000$ events for SPINE (Drielsma et al., 2021b), despite not relying on any labeled examples. This demonstrates that SSL frameworks can robustly capture the complex geometries and energy depositions characteristic of TPC data. However, the suboptimal modeling of the highly-infrequent particles, such as Michel electrons and delta rays, indicates a limitation in our current framework.
Qualitative insights from the model’s internal mechanisms further validate the learned representations and offer an understanding of how these results are achieved. Visualizations of the learned representations via spatially de-biased PCA (i.e., Fig. 8), when compared to the randomly initialized network, show smooth gradients and a clear bias towards individual particle instances. Additionally, visualizations of attention maps from the attention head in the 5th encoder layer (as shown in Figure 9) reveal that PoLAr-MAE learns to selectively focus on voxels belonging to the same particle trajectory. These phenomena are observed for both small, distinct trajectories like Michel electrons and short tracks, as well as for more extended structures such as long tracks and electromagnetic showers. This intrinsic grouping capability is a highly desirable property, and suggests that the learned representations are inherently structured to distinguish between different particle instances. We emphasize that this property is emergent, learned solely from the pretext task of predicting the masked content of LArTPC images from a partial view.
Further architectural innovations are crucial for resolving fine-grained sub-token structures and enhancing model efficiency, for current fixed-resolution, token-based ViT-like models like PoLAr-MAE and Point-MAE can be limited. TPC data is inherently multi-scale with both localized (centimeter-scale) energy depositions and extended (meter-scale) structures. Hierarchical architectures (Wu et al., 2024; Ryali et al., 2023; Fan et al., 2021; Liu et al., 2021; Zhdanov et al., 2025) address the parameter and representational inefficiencies of flat ViT designs, which rely on a fixed resolution, fixed channel approach for all layers throughout the network. By starting spatially fine-grained with fewer channels and becoming spatially coarser yet richer in features, these models enable more efficient parameter use and robust multi-scale feature learning. The success of such hierarchical designs in vision tasks requiring finely detailed per-pixel classification, for example by the Hiera backbone (Ryali et al., 2023) in Meta’s Segment Anything Model 2 (Ravi et al., 2025), clearly underscores their potential. Additionally, point-native architectures, such as Point Transformers (Zhao et al., 2021; Wu et al., 2022, 2024) and Erwin (Zhdanov et al., 2025), offer a promising direction by operating directly on individual point features. We therefore see the exploration of point-native and hierarchical models as a critical next step in advancing LArTPC foundation model research.
Exploring alternative SSL paradigms, such as contrastive learning (Xie et al., 2020; Wilkinson et al., 2025), student-teacher self-distillation for view consistency (Wu et al., 2025) and masked latent prediction (Zeid et al., 2023), local-global view consistency (Wu et al., 2025) may also offer benefits and should be explored in future work.
Clearly, there are many unexplored avenues to pursue in the quest for creating a LArTPC foundation model; in light of this, we hope that the introduction of the PILArNet-M dataset into the community will spur additional research in this domain.
@@
Throughout the development of this work, many changes were introduced to enforce the learning of better representations. We expand on several modifications we tried that did not result in any meaningful change in token semantics, in Appendix F.
@@ Appendix A Dataset
Figure 10: PILArNet-M statistics. Starting from top left and going clockwise: cluster frequencies, i.e., how prevalent each particle type is on a cluster (trajectory) basis; voxel frequencies, i.e., the total semantic ID makeup of all voxels in the dataset; voxel energy distributions, i.e., a histogram of the energies of any occupied voxel across all events in the dataset. The number in the bar plots represents the raw number of each quantity found in the dataset, while the bars themselves are normalized to 1. Note that in the energy deposition distribution, energy values below 0.001 MeV are not present.
The PILArNet-M dataset contains 5.2B points containing energy deposited from 28,251,859 individual particle trajectories, over 1,210,080 individual events. As described Sec. 3 of the main text, each voxel contains information corresponding to not just the energy deposited, but also a fragment ID, group ID, interaction ID, and semantic type. Though we just use the semantic type in this work, we foresee the other IDs being quite useful for other tasks, like, i.e., latent clustering for individual particle (i.e., group ID) identification. Dataset-wide statistics are shown in Figure 10. We also offer visualizations of individual events, where voxels with the same colors correspond to sharing different IDs, in Figure 11. See (Adams et al., 2020) for detailed information about this dataset.
@@
Figure 11: Visual depictions of 7 PILArNet-M events. For each event, voxels are colored according to which integer identifiers they share, for each label type (fragment / cluster ID, group ID, interaction ID, and semantic ID). For ease of visualization, we remove low energy depositions, which would greatly degrade the ability to visualize what each identifier means.
@@ B.1 Comparison of Grouping Strategies
B.1.1 Pareto Frontier of Grouping Methodologies.
@@ 2. Duplicated Points Ratio
Prior work mitigates leakage by using large masking ratios (greater than 60%), but this only reduces overlap probability rather than eliminating it. For particle trajectory datasets—where spatial correlations are strong—even minor overlaps risk exposing the model to trivial reconstruction shortcuts. Our C-NMS strategy directly optimizes the Pareto frontier of $M$ vs. $D$ (see Figure 3 in the main text), ensuring groups are both complete (low missed points) and isolated (low duplication). This mirrors the non-overlapping structure of successful 2D image masking while adapting to the irregular geometry of particle clouds.
Figure 12: Detailed grouping hyperparameter sweep. To find the pareto frontier of grouping methodologies, we sweep over an array of reasonable grouping parameters for each method, reporting the missed ratio $M$ and duplicated ratio $D$ heatmaps for each hyperparameter pair. Clearly, the naïve approaches result in an extreme number of duplicated points in any configuration, while the C-NMS-based grouping method empirically guarantees much smaller amounts of each quantity.
To systematically compare grouping strategies, we perform parameter sweeps for each method. Initial group centers are sampled via FPS (2048 per event). These parameter ranges are described in Table 3. For FPS-based methods, the number of samples per event are defined as $\lceil|\mathcal{X}|/\alpha\rceil$ where $|\mathcal{X}|$ is the number of points in the event. Sweeps generate $80-220$ configurations per method, with Pareto frontiers computed from the missed/duplicated ratio metrics defined in App. B.1.1. Figure 12 provides array of figures showing results for each metric across these sweeps.
@@
B.1.2 Ablation study for pre-training with different grouping strategies
We perform an ablation study on various grouping strategies to quantitatively understand the negative effects that result from using prior grouping methods. We pre-train PoLAr-MAE for 100,000 steps ($\sim$12 epochs) using FPS+$k$-NN, ball query and using C-NMS. We further show ablations on the C-NMS method itself by using extreme overlap factors (0.1, 1.0) and ablating the initial FPS step. We select optimal parameters for each setup using the Pareto frontier study in Appendix B.1.1.
Table 4 provides per-class and overall $F_{1}$ scores from training linear SVMs on the token outputs, as described in Section 3.3 of the main paper. Clearly, using volumetric tokenization greatly increases the effectiveness of the representations learned. Additionally, we also show that it is not necessary to use FPS before C-NMS plus ball query, and not doing so seems to increase representation power by some amount. We do not supply further tests on these pre-trained models on downstream tasks, as we believe the main results in the paper show that the SVM evaluation is generally a good predictor of downstream performance on semantic segmentation fine-tuning.
Table 4: Ablation study on patch grouping strategies. We report the overall $F_{1}$ score and per-class $F_{1}$ on the validation set.
Per-class $F_{1}$
Method Steps $F_{1}$ Track Shower Delta Michel
FPS + kNN 100k 0.469 0.840 0.678 0.339 0.022
FPS + Ball query 100k 0.555 0.943 0.797 0.417 0.061
(Ours) FPS + C-NMS + Ball query 100k 0.687 0.988 0.966 0.449 0.344
$\rightarrow$ Overlap $f{=}0.1$ 100k OOM [^3] OOM: Out of memory.
$\rightarrow$ Overlap $f{=}1.0$ 100k 0.669 0.989 0.967 0.381 0.339
$\rightarrow$ C-NMS + Ball query 100k 0.698 0.989 0.968 0.454 0.383
B.1.3 Runtime comparison with different grouping strategies
Computational efficiency is an important part of foundation model architectures, since low efficiency can considerably increase the pre-training time. We test the efficiency of the C-NMS algorithm compared to prior methods by comparing runtimes of grouping a different number of events. We use the optimal parameters found in the hyperparameter search in App. B.1.1. Fig. 13 shows the relative efficiency of each grouping strategy as a function of batch size. The C-NMS method sits between the efficiencies of the standard $k$-NN and ball query methods, with FPS plus ball query being the most efficient, at $\sim 4,000$ images per second at a batch size of 128. C-NMS plus ball query and without/with the initial FPS step are the second and third most efficient, grouping $\sim 3,300$ and $\sim 2,000$ images per second with the same batch size. Finally, the FPS plus $k$-NN approach is the slowest, at $\sim 1,000$ images per second. Considering the heavy use of $k$-NN based approaches in 3D computer vision literature, we argue the additional computational burden of using C-NMS is small relative to the overall performance of the grouping methods.
Figure 13: Computational efficiency of grouping methods. (Left.) Average time spent to group a single image on the GPU. (Right.) Average number of images able to be grouped per second on the GPU. Our grouping method (C-NMS) is placed between the ball query-based (best) and $k$-NN-based (worst) methods.
@@ Appendix C Hyperparameters
We provide detailed hyperparameter information in the following sections. The full implementation of these models, as well as training code, can be found at https://github.com/DeepLearnPhysics/PoLAr-MAE.
@@ C.1 Grouping strategy
We chose a group radius equal to 5 voxels (15 mm) such that a 1m track will contain  67 mm, and because it is approximately the characteristic distance of trajectory morphologies. To find the best value for the overlap factor $f$, we run a sweep over values of $f$, and examine the missed points ratio $M$ and duplicated points ratio $D$ as defined in Appendix B.1.1. Fig. 14 shows this trade-off for different values of $f$. The value of $f$ that jointly minimizes both these scores is $f=0.73,$ and hence we use this value when pretraining with this grouping strategy.
Figure 14: C-NMS overlap parameter $f$ sweep. We find the optimal $f$ parameter for our group radius of 5 voxels by minimizing both the missed ratio $M$ and duplicated ratio $D$. We find $f=0.73$ to be approximately optimal. Shaded regions represent 1 standard deviation across 32 LArTPC events.
@@ Appendix D Pretraining Results
D.1 Spatial Debiasing of Features
Given a single event, we perform principal component analysis to cast each token’s feature vector to three dimensions. We then normalize the final vectors to $[0,1]^{3}$ and interpret them as RGB channel values. Before PCA, however, we must remove spatial debiasing of individual vectors. Empirically, we found that without this debiasing procedure, tokens were mostly colored as a simple function of their 3-D position. Contrary to usual Vision Transformers, we inject positional encoding at each Transformer layer. This results in embeddings that are rich in spatial context – as needed for the masked modeling pretext task, which heavily relies on the spatial position of masked tokens for reconstruction. However, we are interested in semantic differences between tokens. Thus, we train a simple linear model to predict the position $\textbf{x}=(x,y,z)$ of individual tokens given their feature vector z. We then compute the residual $R$ between the original and predicted feature vectors, and finally perform PCA.
Concretely, let the dataset
$\mathcal{D}=\bigl{\{}(\mathbf{x}_{i},\mathbf{z}_{i})\bigr{\}}_{i=1}^{N},\qquad\mathbf{x}_{i}\in\mathbb{R}^{3},\;\mathbf{z}_{i}\in\mathbb{R}^{D_{\text{embed}}},\;D_{\text{embed}}=384.$
We model each latent vector as an affine function of its 3‑D position:
$\hat{\mathbf{z}}_{i}=W\mathbf{x}_{i}+\mathbf{b},$
where $W\in\mathbb{R}^{D_{\text{embed}}\times 3}$ and $\mathbf{b}\in\mathbb{R}^{D_{\text{embed}}}$ are learned parameters. The optimal $(W^{\star},\mathbf{b}^{\star})$ are obtained by ordinary least squares:
$(W^{\star},\mathbf{b}^{\star})=\arg\min_{W,\mathbf{b}}\sum_{i=1}^{N}\bigl{\lVert}\mathbf{z}_{i}-W\mathbf{x}_{i}-\mathbf{b}\bigr{\rVert}_{2}^{2}.$
In matrix form, we augment the position matrix with a bias column:
$X=\begin{bmatrix}\mathbf{x}_{1}^{\!\top}\\[-2.0pt] \vdots\\ \mathbf{x}_{N}^{\!\top}\end{bmatrix}\in\mathbb{R}^{N\times 3},\quad\tilde{X}=\begin{bmatrix}X&\mathbf{1}\end{bmatrix}\in\mathbb{R}^{N\times 4},\quad Z=\begin{bmatrix}\mathbf{z}_{1}^{\!\top}\\[-2.0pt] \vdots\\ \mathbf{z}_{N}^{\!\top}\end{bmatrix}\in\mathbb{R}^{N\times D_{\text{embed}}}.$
Let $\tilde{W}=\begin{bmatrix}W&\mathbf{b}\end{bmatrix}\in\mathbb{R}^{D_{\text{embed}}\times 4}$. The closed‑form solution is
$\tilde{W}^{\star}=\bigl{(}\tilde{X}^{\top}\tilde{X}\bigr{)}^{-1}Z^{\top}\tilde{X}.$
For each sample, subtract the fitted component to obtain the residual
$\mathbf{r}_{i}=\mathbf{z}_{i}-\hat{\mathbf{z}}_{i}=\mathbf{z}_{i}-W^{\star}\mathbf{x}_{i}-\mathbf{b}^{\star}.$
Collecting all residuals,
$R=Z-\tilde{X}\tilde{W}^{\star\top}\in\mathbb{R}^{N\times D_{\text{embed}}},$
which can now be used for downstream analysis without linear spatial bias.
D.2 Extra PCA Results
In Figure 16, we show additional visualizations of representations learned by both Point-MAE and PoLAr-MAE.
@@ Appendix E Segmentation Results
In Figure 17, we show additional examples of semantic segmentation results for both Point-MAE and PoLAr-MAE fine-tuned models. In Table 7, we provide additional evaluation metrics for the fine-tuned Point-MAE and PoLAr-MAE models. We also supply confusion matrices for each model, normalized across predictions, in Figure 18.
@@
Figure 17: Further examples of dense classification. Model types are abbreviated with PL meaning PoLAr-MAE, PM meaning Point-MAE, PEFT meaning parameter-efficient fine-tuning, and FFT meaning full fine-tuning. Best viewed zoomed in.
@@
Point-MAE PEFT 0.772 0.965 0.983 0.569 0.572 0.702 0.990 0.998 0.413 0.406 0.948 0.941 0.968 0.919 0.967
Point-MAE FFT 0.831 0.963 0.994 0.561 0.807 0.771 0.996 0.997 0.395 0.697 0.962 0.932 0.990 0.969 0.957
@@
Figure 18: Semantic segmentation confusion matrices. Rows denote fine-tuning strategies: PEFT and FFT. Columns denote the pre-trained model used: Point-MAE and PoLAr-MAE. Each column sums to 1, and the diagonal values represent precision metrics for each class.